OpenStudy (anonymous):

a cars engine has 6 cylinders, each have a diameter of 9.6 cm and height of 10.4 cm. Find the capacity of the engines cylinders in cubic cm? Help me PLEASE!! :\

OpenStudy (diyadiya):

\[\Large Volume \ of \ a \ cylinder = \pi \times (radius)^ 2 \times height \]

OpenStudy (anonymous):

I have that part but how do i change the answer 752.39 into Cubic Centimeters?

OpenStudy (diyadiya):

So you got the capacity as 752.39 ? then its in cubic cm itself 759.39 cm^3 (cubic cm)

OpenStudy (anonymous):

yes i did 3.14x23.04x10.4 which gave me the answer 752.39, so i just add the 3 to the end and make it cubed? Correct?

OpenStudy (diyadiya):

No its not like adding 3 to the end unit of volume is \(\Large cm^3 \) If the radius & height was in meters then the volume would have been \(\Large m^3\) so capacity of one cylinder is \(\Large 752.39 cm^3 \)

OpenStudy (anonymous):

the radius and height were in CM also

OpenStudy (diyadiya):

yeah since they are in cm we have the unit of volume as \(\Large cm^3\)

OpenStudy (anonymous):

okay, thanks, that helped me out alot! thank you soooooooooo much :)

OpenStudy (diyadiya):

Anytime =)